[figure: 9.7 cm]

MAGNETIC PROPERTIES OF SPIN-LADDERS (A_{2}Cu_{2}O_{3})_{m}(CuO_{2})_{n}

M. Baran^{a}, H. Szymczak^{a}, R. Szymczak^{a}, P. Aleshkevych^{a}, S. Dyeyev^{a},

L. Leonyuk^{b}, and G.-J. Babonas^{c} a Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, PL-02-668 Warsaw, Poland b Moscow State University, Moscow, Russia c Semiconductor Physics Institute, 2600 Vilnius, Lithuania

Magnetization measurements have been performed on several single crystals of the S=1/2 quasi one dimensional system (A_{2}Cu_{2}O_{3})_{m}(CuO_{2})_{n} (A=Sr, Ca, Y, Nd, Pr + some small admixtures) in temperature range 5 K - 300 K and in wide magnetic field range. The studied crystals with m/n = 5/7, 7/10 and 9/13, of the orthorhombic structure, are incommensurate along the c-axis and in a commensurate approximation are characterized by approximate lattice constants a=11.3 Å, b=12.8 Å and c=19.5 Å, 27.5 Å, and 35.6 Å, respectively. In the crystal structure of investigated compounds there can be distinguished both simple CuO_{2} chains and two-leg spin ladders of copper ions. The analysis of the temperature dependences of magnetic susceptibility indicates that the chain in these compounds has a spin gap in the excitation spectrum. The origin of this spin gap is yet unclear. The values of this spin gap are from about 30 K to 130 K for crystals with various compositions and structures. The contribution to the susceptibility both of Cu^{2+} ions located in various sites as well as other ions located in A-site has been estimated.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Baran

Address for correspondence : Dr M. Baran Al. Lotników 32/46 02-668 Warszawa

Email address : baran@ifpan.edu.pl